# OG-VLA: 3D-Aware Vision Language Action Model via Orthographic Image Generation

### v1

@@ Abstract
We introduce OG-VLA, a novel architecture and learning framework that combines the generalization strengths of Vision Language Action models (VLAs) with the robustness of 3D-aware policies. We address the challenge of mapping natural language instructions and multi-view RGBD observations to quasi-static robot actions. 3D-aware robot policies achieve state-of-the-art performance on precise robot manipulation tasks, but struggle with generalization to unseen instructions, scenes, and objects. On the other hand, VLAs excel at generalizing across instructions and scenes, but can be sensitive to camera and robot pose variations. We leverage prior knowledge embedded in language and vision foundation models to improve generalization of 3D-aware keyframe policies. OG-VLA projects input observations from diverse views into a point cloud which is then rendered from canonical orthographic views, ensuring input view invariance and consistency between input and output spaces. These canonical views are processed with a vision backbone, a Large Language Model (LLM), and an image diffusion model to generate images that encode the next position and orientation of the end-effector on the input scene. Evaluations on the Arnold and Colosseum benchmarks demonstrate state-of-the-art generalization to unseen environments, with over 40% relative improvements while maintaining robust performance in seen settings. We also show real-world adaption in 3 to 5 demonstrations along with strong generalization. Videos and resources at https://og-vla.github.io
1 Introduction
@@
We study the problem of mapping natural language instructions and multiple posed RGBD observations to robot actions, with specific focus on quasi-static manipulation tasks that can be decomposed into a sequence of end-effector keyframes. This category encompasses a wide variety of tasks such as pick-and-place, opening/closing doors and containers, manipulating buttons, valves, switches, and more. Building robust policies that solve such tasks in unseen environments remains an open challenge that could enable numerous industrial and household applications, from cleaning and sorting robots to machine tending.
VLAs have recently demonstrated successful generalization to concepts unseen in the robotics training data [1, 2, 3, 4], such as manipulating novel objects based on language instructions. While achieving generalization breakthroughs, they require massive training datasets [3, 5, 4] and typically accept a single RGB view input. As a result, despite the large amount of training data, the resulting systems remain sensitive to variations in camera and robot poses, which hurt their adaptability to new applications [6]. They also lack explicit visual reasoning when predicting actions (often as LLM tokens), constraining their ability to perform precise, generalizable 3D spatial reasoning.
In contrast, 3D-aware keyframe based policies learn effectively from few demonstrations and generalize well to novel camera poses and object placements [7, 8, 9, 10] as confirmed in the camera perturbation evaluation study by Pumacay et al. [11]. This success stems from a 3D scene representation within the model, such as a voxel map [7], canonical orthographic views [8], or point clouds [10]. However, unlike VLAs, these systems overfit to the training scenes and objects, failing to accept instructions that refer to new, previously unseen objects.
We propose OG-VLA (Orthographic-image Generation Vision-Language-Action model), a novel robot policy architecture that combines the generalization strengths of VLAs with robustness of 3D-aware policies. OG-VLA uses LLMs and image generation to map posed RGBD images and language to 6-DOF end-effector pose keyframes in a sequence. The system comprises four components: a point cloud renderer that renders a scene reconstruction to canonical orthographic views, a vision backbone that encodes these views into visual embeddings, an LLM that predicts action tokens, and an image diffusion model that decodes these action tokens to predict actions on each of the orthographic views through image generation, which we decode to final 3D poses. The LLM and image diffusion models are trained end-to-end, so that they work together to produce consistent and precise predictions required for robotic manipulation. Figure 1 illustrates our method.
@@
We conduct simulation and real-robot experiments. In simulation, the Arnold [12] benchmark tests generalization to unseen objects and environments, while Colosseum [11] evaluates robustness to variations in camera poses, object poses, colors, and distractors. We show significantly improved performance on generalization tests of both benchmarks, achieving state-of-the-art on Arnold [12]. In real-world experiments, we show our method’s ability to learn manipulation tasks from as few as 3 to 5 demonstrations, highlighting its suitability for kinesthetic teaching and rapid adaptation to new domains. We also present a detailed study of our model architecture design choices, providing insights into its functioning as well as potential future improvements.
2 Related Work
@@
Large Language Models have seen an explosion in research on their use-cases in robotics, such as task planning [13, 14, 15], reward generation for reinforcement learning [16, 17], and interfacing with vision models [18]. However, although these methods generalize at a high level, they assumed access to low-level skills such pick, place, open, and close, along with perception and scene representation systems. These skills are hard to build, and their development is in line with the goals of this work.
@@
3D-aware keyframe-based multi-task policies have shown the ability to learn complex manipulation behaviors from as little as ten demonstrations per task. Examples of these works include PerAct [7] based on voxel grids, RVT [8, 9] based on orthonormal views, and Act3D [10] based visual feature point clouds as the scene representation. These systems significantly improve upon image-based policies [19, 20] in the amount of data they require and their robustness to new object placements at test-time. However, these systems have been trained from scratch on specific tasks, robots, and environments, and struggle to generalize to new objects and scenes, or instructions.
@@
Vision Language Action models (VLAs) leverage large-scale prior knowledge from LLMs and vision foundation models to create robot policies that generalize to new concepts and objects at test time [1, 21, 2, 3, 22, 4]. However, these systems require huge amount of demonstration data to train. For example, Kim et al. [3] use over 900k demos from the Open-X Embodiment dataset [23]. Despite the large quantity of data, the resulting systems are sensitive to changes in the 3D environment, such as different camera poses relative to the robot system. Our system trains with significantly less data to achieve state-of-the-art performance and generalization.
@@
Generative image and video models have been explored for robot policies as well. Du et al. [24] explore mapping generated videos of robots back to control via inverse dynamics. Genima [25] makes this easier by drawing robot joint annotations as textured spheres on the video, while RT-Trajectory [26] generates trajectory annotations on images. In contrast, we predict annotations on 3D canonical views, which allows us to more easily solve for 3D end-effector poses even in free space, and improves generalization from few demonstrations by enabling SE(3) data augmentation. We show that our model can do free space reasoning for tasks such as lift the bottle 30cm off the ground (Figure 5). Methods that directly predict heatmaps on images have also been previously used for language-guided navigation tasks [27, 28].
@@
3 Method: Orthographic-image Generation Vision-Language-Action model
@@
At deployment time, the input to our system is a language instruction $l$, and a set of observations $O_{k}=\{I_{k},D_{k},P_{k},K_{k}\}$, where $I_{k}$ is an RGB image, $D_{k}$ is a corresponding depth image, $P_{k}$ is the camera pose, and $K_{k}$ are the camera intrinsics, with a camera index $k$. The output of our system is an end-effector state $s=\langle p,\omega\rangle$, which consists of a position target $p$, rotation target $\omega$. To complete a task, we sequentially execute our system, at each time-step using a motion planner to reach the predicted $s$, and obtain the next set of observations. Figure 2 shows our model architecture.
3.1 Multi-Modal Vision and Language Model
@@
At the core of our system is a Large Language Model (LLM). The LLM takes as input a sequence of input tokens (vectors) $\langle t_{1},\dots,t_{i}\rangle$, and generates a sequence of output tokens (vectors) $\langle t_{i+1},\dots,t_{i+j}\rangle$. We use three types of input and output tokens: (1) text tokens, computed from a text tokenizer and embedding table, (2) input image tokens, which are either patch tokens or the image CLS token [29], computed by a visual encoder, and projected to LLM space through a learned MLP input projection, and (3) output image (action) tokens, which we add to the LLM vocabulary and decode as a special token using an additional MLP decoder. The output image tokens represent the next robot action. We use an image diffusion model to decode the image tokens into actions by producing images that contain annotations that illustrate the gripper position and rotation over a set of input views of the scene. The end-effector state $s$ is decoded from these image annotations.
@@
3.2 3D-Aware Reasoning with Orthogonal Orthographic Projections
To imbue the LLM with 3D-awareness, we unproject all input camera images into a point cloud in a canonical workspace. We then render the scene from a fixed set of views (independent of the input camera poses) before feeding them to the LLM. This brings the input and output in the same space, and our selection of the views (orthogonal views such as "front", "top", "left, "right", rendered in orthographic mode) ensures no ambiguity in output.
Input Reprojection to Canonical Views. For each camera observation $\{I_{k},D_{k},P_{k},K_{k}\}$ with $N_{k}$ valid depth pixels, we compute a point cloud $C_{k}\in\mathbbm{R}^{N_{k}\times 6}$, where each contains the RGB color and 3D coordinate in a fixed reference frame. We compute an aggregated point cloud $C=\bigcup_{k=1}^{K}C_{k}$ overall all input cameras. Next, we define a set of $m$ canonical cameras $\{P^{C}_{c},K^{C}_{c}\}_{c=1,\dots,m}$, where each $P^{C}_{c}$ is a camera pose and $K^{C}_{c}$ are the intrinsics for a given orthograpic camera $c$. We then render the point cloud to an RGB image $I^{C}_{c}\in\mathbbm{R}^{h\times w\times 3}$ seen by each orthographic camera.
Our system allows adapting the set of canonical camera parameters based on application and workspace geometry. In this work, we use four orthographic cameras that view the workspace from front, left, right and top directions, such that the workspace fills the camera image. Figure 2 shows the input RGBD views and the resulting orthographic images. While we use point clouds, our method may accommodate any 3D representation (e.g., neural radiance fields [30, 31] or novel-view synthesis methods [32]) that supports canonical view rendering.
@@
We feed the input sequence to the LLM to produce an output sequence of format: $\langle t^{1}_{a},t^{2}_{a},t^{3}_{a},t^{4}_{a},t^{o}_{1},t^{o}_{2},\dots,t^{o}_{j}\rangle,$ where $t^{(\cdot)}_{a}$ are four image tokens that together represent the next action, and $t^{o}_{(\cdot)}$ are accompanying text tokens as a response to the input prompt. We show the abbreviated prompt in Figure 2. Full prompt and text response is provided in the Appendix B.
@@
where $H_{c}\in\{\mathbbm{R}^{h,w,3}\}$ is a reconstruction of the input canonical image $I^{C}_{c}$ with overlaid annotations that encode the gripper position and orientation. Each image is aligned with one of the $m$ canonical input views $c$, as shown in Figure 2. While any image generator can be used, we use StableDiffusion 1.5 [33]. $e^{i}_{a}$ acts as the textual conditioning and $e_{c}^{x}$ as the visual conditioning for the ImageGenerator. In practice, we use $e^{i}_{a}+\text{CLIP}(\textsc{Prompt}(l))$ for textual conditioning, essentially adding a residual skip connection from a direct text encoding, a design choice informed by empirical experiments. CLIP is the textual encoder used for pretraining the StableDiffusion 1.5 model.
@@
$p^{hm}=\arg\max_{p}\prod_{c=1,\dots,m}(H_{c}[\textsc{CameraProjection}(p,P^{C}_{c},K^{C}_{c})]+\epsilon)\;,$ (2)
where CameraProjection projects the 3D point $p$ to 2D image coordinates w.r.t. the orthographic camera $c$. The square brackets represent a 2D pixel-wise indexing operation with interpolation to support sub-pixel coordinates. $\epsilon$ is a small value we add to allow decoding in situations where one of the heatmaps is zero for all 3D points.
@@
3.3 Training and Implementation Details
Our architecture builds on X-VILA [34], a multi-modal chat model supporting language, visual, video, and audio modalities. OG-VLA is initialized from X-VILA’s pretrained weights to leverage its any-to-any modality alignment, focusing on text-image input to text-image output alignment. We leave study of enriching human-robot interaction using the other modalities to future work. We train OG-VLA using DeepSpeed [35]. We freeze the visual encoder and tune the LLM, input and output projection networks, and ImageGenerator with end-to-end gradient flow. Following X-VILA, we use ImageBind [36] as the visual encoder, linear layers for input and output projections, and Stable Diffusion 1.5 [37] for the ImageGenerator. The LLM is based on Vicuna-7b v1.5 architecture. ImageBind encodes images into 256 patches (16×16) with a CLS token. Each training sample includes a natural language instruction $l$, visual observations ${I_{k},D_{k},P_{k},K_{k}}$, and ground-truth gripper state $\hat{s}$. Following prior work [7, 8], we augment each keyframe with $N$ SE(3)-transformed purterbations: translation in [±0.1m, ±0.1m, ±0.1m], rotation in [±0°, ±0°, ±90°]. All models are trained on 8×A100 GPUs with batch size 64. All models are trained for one run, and the LLM is LoRA-finetuned. Inference is done on a single A100 GPU. For inference-time image sampling from Stable Diffusion 1.5, we use 100 steps and a guidance scale of 7.0. These parameters result in reasonable sampled image quality and avoid jitters in the sampled image from the model’s latent space.
4 Simulation Experiments

### v2

# OG-VLA: Orthographic Image Generation for 3D-Aware Vision-Language Action Model

@@ Abstract
We introduce OG-VLA, a novel architecture and learning framework that combines the generalization strengths of Vision Language Action models (VLAs) with the robustness of 3D-aware policies. We address the challenge of mapping natural language instructions and one or more RGBD observations to quasi-static robot actions. 3D-aware robot policies achieve state-of-the-art performance on precise robot manipulation tasks, but struggle with generalization to unseen instructions, scenes, and objects. On the other hand, VLAs excel at generalizing across instructions and scenes, but can be sensitive to camera and robot pose variations. We leverage prior knowledge embedded in language and vision foundation models to improve generalization of 3D-aware keyframe policies. OG-VLA unprojects input observations from diverse views into a point cloud which is then rendered from canonical orthographic views, ensuring input view invariance and consistency between input and output spaces. These canonical views are processed with a vision backbone, a Large Language Model (LLM), and an image diffusion model to generate images that encode the next position and orientation of the end-effector on the input scene. Evaluations on the Arnold and Colosseum benchmarks demonstrate state-of-the-art generalization to unseen environments, with over 40% relative improvements while maintaining robust performance in seen settings. We also show real-world adaption in 3 to 5 demonstrations along with strong generalization. Videos, additional details, and resources at https://og-vla.github.io
I Introduction
We study the problem of mapping natural language instructions and one or more posed RGBD observations to robot actions, with specific focus on quasi-static manipulation tasks that can be decomposed into a sequence of end-effector keyframes. This category encompasses a wide variety of tasks such as pick-and-place, opening/closing doors and containers, manipulating buttons, valves, switches, and more. Building robust policies that solve such tasks in unseen environments remains an open challenge that could enable numerous industrial and household applications, from cleaning and sorting robots to machine tending.
VLAs have recently demonstrated successful generalization to concepts unseen in the robotics training data [driess2023palm, brohan2023rt2, kim24openvla, pertsch2025pi0fast], such as manipulating novel objects based on language instructions. While achieving generalization breakthroughs, they require massive training datasets [kim24openvla, openxembodiment, pertsch2025pi0fast] and typically accept a single RGB view input. As a result, despite the large amount of training data, the resulting systems remain sensitive to variations in camera and robot poses, which hurt their adaptability to new applications [hamster2025]. They also lack explicit visual reasoning when predicting actions (often as LLM tokens), constraining their ability to perform precise, generalizable 3D spatial reasoning.
In contrast, 3D-aware keyframe based policies learn effectively from few demonstrations and generalize well to novel camera poses and object placements [shridhar2022peract, goyal2023rvt, goyal2024rvt, gervet2023act3d] as confirmed in the camera perturbation evaluation study in [2024colosseum]. This success stems from a 3D scene representation within the model, such as a voxel map [shridhar2022peract], canonical orthographic views [goyal2023rvt], or point clouds [gervet2023act3d]. However, unlike VLAs, these systems overfit to the training scenes and objects, failing to accept instructions that refer to new, previously unseen objects.
We propose OG-VLA (Orthographic-image Generation Vision-Language-Action model), a novel robot policy architecture that combines the generalization strengths of VLAs with robustness of 3D-aware policies. OG-VLA uses LLMs and image generation to map one or more posed RGBD images and language to 6-DOF end-effector pose keyframes in a sequence. The system comprises four components: a point cloud renderer that renders a scene reconstruction to canonical orthographic views, a vision backbone that encodes these views into visual embeddings, an LLM that predicts action tokens, and an image diffusion model that decodes these action tokens to predict actions on each of the orthographic views through image generation, which we decode to final 3D poses. The LLM and image diffusion models are trained end-to-end, so that they work together to produce consistent and precise predictions required for robotic manipulation. Figure 1 illustrates our method.
@@ I Introduction
We conduct simulation and real-robot experiments. In simulation, the Arnold [gong2023arnold] benchmark tests generalization to unseen objects and environments, while Colosseum [2024colosseum] evaluates robustness to variations in camera poses, object poses, colors, and distractors. We show significantly improved performance on generalization tests of both benchmarks, achieving state-of-the-art on Arnold [gong2023arnold]. In real-world experiments, we show our method’s ability to learn manipulation tasks from as few as 3 to 5 demonstrations, highlighting its suitability for kinesthetic teaching and rapid adaptation to new domains. We also present a detailed study of our model architecture design choices, providing insights into its functioning as well as potential future improvements.
II Related Work
@@
Large Language Models have seen an explosion in research on their use-cases in robotics, such as task planning [saycan2022arxiv, singh2023progprompt, codeaspolicies2022], reward generation for reinforcement learning [ma2023eureka, yu2023language], and interfacing with vision models [huang2023voxposer]. However, although these methods generalize at a high level, they assumed access to low-level skills such pick, place, open, and close, along with perception and scene representation systems. These skills are hard to build, and their development is in line with the goals of this work.
@@
3D-aware keyframe-based multi-task policies have shown the ability to learn complex manipulation behaviors from as little as ten demonstrations per task. Examples of these works include PerAct [shridhar2022peract] based on voxel grids, RVT [goyal2023rvt, goyal2024rvt] based on orthonormal views, and Act3D [gervet2023act3d] based visual feature point clouds as the scene representation. These systems significantly improve upon image-based policies [r3m, mvp] in the amount of data they require and their robustness to new object placements at test-time. However, these systems have been trained from scratch on specific tasks, robots, and environments, and struggle to generalize to new objects and scenes, or instructions.
@@
Vision Language Action models (VLAs) leverage large-scale prior knowledge from LLMs and vision foundation models to create robot policies that generalize to new concepts and objects at test time [driess2023palm, brohan2023rt2, kim24openvla, black2024pi0, pertsch2025pi0fast, intelligence2025pi05visionlanguageactionmodelopenworld]. However, these systems require huge amount of demonstration data to train. For example, [kim24openvla] uses over 900k demos from the Open-X Embodiment dataset [padalkar2023open]. Despite the large quantity of data, the resulting systems are sensitive to changes in the 3D environment, such as different camera poses relative to the robot system. Our system trains with significantly less data to achieve state-of-the-art performance and generalization.
@@
Generative image and video models have been explored for robot policies as well. [du2023videogenpolicies] explores mapping generated videos of robots back to control via inverse dynamics. Genima [shridhar2024generative] makes this easier by drawing robot joint annotations as textured spheres on the video, while RT-Trajectory [gu2023rt] generates trajectory annotations on images. In contrast, we predict annotations on 3D canonical views, which allows us to more easily solve for 3D end-effector poses even in free space, and improves generalization from few demonstrations by enabling SE(3) data augmentation. We show that our model can do free space reasoning for tasks such as lift the bottle 20 cm off the ground (Figure 2). Methods that directly predict heatmaps on images have also been previously used for language-guided navigation tasks [anderson2019chasing, blukis2019learning].
@@
III Method: Orthographic-image Generation Vision-Language-Action model
@@
At deployment time, the input to our system is a language instruction $l$ and a set of observations $O_{k}=\{I_{k},D_{k},P_{k},K_{k}\}$, where $I_{k}$ is an RGB image, $D_{k}$ is a corresponding depth image, $P_{k}$ is the camera pose, and $K_{k}$ are the camera intrinsics, with a camera index $k$. The output of our system is an end-effector state $s=\langle p,\omega\rangle$, which consists of a position target $p$, rotation target $\omega$. To complete a task, we sequentially execute our system, at each time-step using a motion planner to reach the predicted $s$, and obtain the next set of observations. Figure 2 shows our model architecture.
III-A Multi-Modal Vision and Language Model
@@
At the core of our system is a Large Language Model (LLM). The LLM takes as input a sequence of input tokens (vectors) $\langle t_{1},\dots,t_{i}\rangle$, and generates a sequence of output tokens (vectors) $\langle t_{i+1},\dots,t_{i+j}\rangle$. We use three types of input and output tokens: (1) text tokens, computed from a text tokenizer and embedding table, (2) input image tokens, which are either patch tokens or the image CLS token [devlin2018bert], computed by a visual encoder, and projected to LLM space through a learned MLP input projection, and (3) output image (action) tokens, which we add to the LLM vocabulary and decode as a special token using an additional MLP decoder. The output image tokens represent the next robot action. We use an image diffusion model to decode the image tokens into actions by producing images that contain annotations that illustrate the gripper position and rotation over a set of input views of the scene. The end-effector state $s$ is decoded from these image annotations.
@@
III-B 3D-Aware Reasoning with Orthogonal Orthographic Projections
To imbue the LLM with 3D-awareness, we unproject all input camera images into a point cloud in a canonical workspace. We then render the scene from a fixed set of views (independent of the input camera poses) before feeding them to the LLM. This brings the input and output in the same space, and our selection of the views (orthogonal views such as ‘front’, ‘top’, ‘left’, ‘right’, rendered in orthographic mode) ensures no ambiguity in output.
Input Reprojection to Canonical Views. For each camera observation $\{I_{k},D_{k},P_{k},K_{k}\}$ with $N_{k}$ valid depth pixels, we compute a point cloud $C_{k}\in\mathbbm{R}^{N_{k}\times 6}$, where each contains the RGB color and 3D coordinate in a fixed reference frame. We compute an aggregated point cloud $C=\bigcup_{k=1}^{K}C_{k}$ over all input cameras. Next, we define a set of $m$ canonical cameras $\{P^{C}_{c},K^{C}_{c}\}_{c=1,\dots,m}$, where each $P^{C}_{c}$ is a camera pose and $K^{C}_{c}$ are the intrinsics for a given orthograpic camera $c$. We then render the point cloud to an RGB image $I^{C}_{c}\in\mathbbm{R}^{h\times w\times 3}$ seen by each orthographic camera.
Our system allows adapting the set of canonical camera parameters based on application and workspace geometry. In this work, we use four orthographic cameras that view the workspace from front, left, right and top directions, such that the workspace fills the camera image. Figure 2 shows the input RGBD views and the resulting orthographic images. While we use point clouds, our method may accommodate any 3D representation (e.g., neural radiance fields [mildenhall2021nerf, yu2021pixelnerf] or novel-view synthesis methods [kulhanek2022viewformer]) that supports canonical view rendering.
@@
We feed the input sequence to the LLM to produce an output sequence of format: $\langle t^{1}_{a},t^{2}_{a},t^{3}_{a},t^{4}_{a},t^{o}_{1},t^{o}_{2},\dots,t^{o}_{j}\rangle,$ where $t^{(\cdot)}_{a}$ are four image tokens that together represent the next action, and $t^{o}_{(\cdot)}$ are accompanying text tokens as a response to the input prompt. We show the abbreviated prompt in Figure 2. Full prompt and text response is provided in the Appendix on project website.
@@
where $H_{c}\in\{\mathbbm{R}^{h,w,3}\}$ is a reconstruction of the input canonical image $I^{C}_{c}$ with overlaid annotations that encode the gripper position and orientation. Each image is aligned with one of the $m$ canonical input views $c$, as shown in Figure 2. While any image generator can be used, we use StableDiffusion 1.5 [rombach2021stablediffusion]. $e^{i}_{a}$ acts as the textual conditioning and $e_{c}^{x}$ as the visual conditioning for the ImageGenerator. In practice, we use $e^{i}_{a}+\text{CLIP}(\textsc{Prompt}(l))$ for textual conditioning, essentially adding a residual skip connection from a direct text encoding, a design choice informed by empirical experiments. CLIP is the textual encoder used for pretraining the StableDiffusion 1.5 model.
@@
$p^{hm}=\operatorname*{arg\,max}_{p}\prod_{c=1,\dots,m}(H_{c}[\mathcal{C}(p,P^{C}_{c},K^{C}_{c})]+\epsilon)$ (2)
where $\mathcal{C}$ projects the 3D point $p$ to 2D image coordinates w.r.t. the orthographic camera $c$. The square brackets represent a 2D pixel-wise indexing operation with interpolation to support sub-pixel coordinates. $\epsilon$ is a small value we add to allow decoding in situations where one of the heatmaps is zero for all 3D points.
@@
III-C Training and Implementation Details
Our architecture builds on X-VILA [ye2024xvila], a multi-modal chat model supporting language, visual, video, and audio modalities. OG-VLA is initialized from X-VILA’s pretrained weights to leverage its any-to-any modality alignment, focusing on text-image input to text-image output alignment. We leave study of enriching human-robot interaction using the other modalities to future work. We train OG-VLA using DeepSpeed [rasley2020deepspeed]. We freeze the visual encoder and tune the LLM, input and output projection networks, and ImageGenerator with end-to-end gradient flow. Following X-VILA, we use ImageBind [girdhar2023imagebind] as the visual encoder, linear layers for input and output projections, and Stable Diffusion 1.5 [sd15] for the ImageGenerator. The LLM is based on Vicuna-7b v1.5 architecture. ImageBind encodes images into 256 patches (16×16) with a CLS token. Each training sample includes a natural language instruction $l$, visual observations $I_{k},D_{k},P_{k},K_{k}$}, and ground-truth gripper state $\hat{s}$. Following prior work [shridhar2022peract, goyal2023rvt], we augment each keyframe with $N$ SE(3)-transformed purterbations: translation in [±0.1m, ±0.1m, ±0.1m], rotation in [±0°, ±0°, ±90°]. All models are trained on 8×A100 GPUs with batch size 64. All models are trained for one run, and the LLM is LoRA-finetuned. Inference is done on a single A100 GPU. For inference-time image sampling from Stable Diffusion 1.5, we use 100 steps and a guidance scale of 7.0. These parameters result in reasonable sampled image quality and avoid jitters in the sampled image from the model’s latent space.
IV Simulation Experiments